# Minimal Thermodynamic Cost of Computing with Circuits

### v2

@@ Appendix B Proofs
The mismatch cost associated with the execution of a gate $\mu$ in the middle of a run is given by Eq. (24),
$\mathcal{MC}_{\mu}=\mathrm{I}_{0}(\boldsymbol{x}_{\mu};\boldsymbol{x}_{\mathrm{ch}(\mu)})+\Delta_{\mu}$ (58)
@@
$\Delta_{\mu}:=D(p_{0}(\boldsymbol{x}_{\mu},\boldsymbol{x}_{\mathrm{pa}(\mu)})||q_{0}(\boldsymbol{x}_{\mu},\boldsymbol{x}_{\mathrm{pa}(\mu)}))-D(p_{0}(\boldsymbol{x}_{\mathrm{pa}(\mu)})||q_{0}(\boldsymbol{x}_{\mathrm{pa}(\mu)})).$
Note that $p_{0}(\boldsymbol{x}_{\mu},\boldsymbol{x}_{\mathrm{pa}(\mu)})$ is the distribution over the joint state of $\mu$ and $\mathrm{pa}(\mu)$ before the execution of gate $\mu$ when $\mu$ and $\mathrm{pa}(\mu)$ are independent of each other. Therefore, $p_{0}(\boldsymbol{x}_{\mu},\boldsymbol{x}_{\mathrm{pa}(\mu)})$ can be written as a product distribution,
$p_{0}(\boldsymbol{x}_{\mu},\boldsymbol{x}_{\mathrm{pa}(\mu)})=p_{0}(\boldsymbol{x}_{\mu})p_{0}(\boldsymbol{x}_{\mathrm{pa}(\mu)})$ (59)
For simplification we assume that $q_{0}$ is a product distribution, i.e., $q_{0}(\boldsymbol{x}_{\mu},\boldsymbol{x}_{\mathrm{pa}(\mu)})=q_{0}(\boldsymbol{x}_{\mu})q_{0}(\boldsymbol{x}_{\mathrm{pa}(\mu)})$. In that case, Eq. 58 simplfies to
$\mathcal{MC}_{\mu}=\mathrm{I}_{\mu}(\mathrm{X}_{\mu};\mathrm{X}_{\mathrm{ch}(\mu)})+D\left(p_{\mu}||q_{\mu}\right).$ (60)
The mutual information between a gate $\mu$ and its children gates $\mathrm{ch}(\mu)$ is always upper bounded by the entropy of the gate $\mu$,
$\mathcal{MC}_{\mu}$ $\leq S(p_{\mu})+D(p_{\mu})\|q_{\mu}))$ (61)
$=\mathcal{C}~{}(p_{\mu},q_{\mu})$ (62)
where $\mathcal{C}~{}(p_{\mu},q_{\mu})$ is the cross entropy of the distribution $p_{\mu}$ relative to a distribution $q_{\mu}$. Moreover, the above inequality is an equality when $\boldsymbol{x}_{\mu}$ is completely determined by $\boldsymbol{x}_{\mathrm{ch}(\mu)}$ ($S(\mathrm{X}_{\mu}|\mathrm{X}_{\mathrm{ch}(\mu)})=0$).
Furthermore, regardless of the actual distribution $p_{\mu}(\boldsymbol{x}_{\mu})$, the cross-entropy is upper bounded by $\ln(1/q_{\mu}^{\min})$. Therefore,
$\mathcal{MC}_{\mu}\leq\ln\left(\frac{1}{q_{\mu}^{\min}}\right)$ (63)
If all the gates in the circuit have the same prior, the sum total of the mismatch cost of running all the gates is upper bounded by,
$\sum_{\mu\in V_{nin}}\mathcal{MC}_{max}=|C_{n}|\ln\left(\frac{1}{q_{0}^{\min}}\right)$ (64)
where $|C_{n}|$ is the number of gates in the circuit. Note that the Eq. 64 does not include the cost of overwriting. So the true total cost is going to be the sum of (64) and the overwriting cost. From Eq. (20), the mismatch cost associated with the overwriting is upper bounded by $S(p_{in})$. Therefore, total associated mismatch cost of a circuit is,
$\mathcal{MC}(C_{n})\leq|C_{n}|\ln\left(\frac{1}{q^{\min}}\right)+S(p_{in})$ (65)
@@
From Eq. 63, the mismatch cost for each gate is upper bound by $\ln(1\backslash q_{\mu}^{\min})$. Let us denote thee prior distribution of logic gate $g\in\mathcal{B}$ as $q_{g}$, and $K_{g}=\ln(1\backslash q_{\mu}^{\min})$. If #$\#_{g}(n)$ is the number of gates of type $g\in\mathcal{B}$ in a circuit $C_{n}$, then the sum total of mismatch cost of gates is upper bounded.
$\sum_{\mu=0}^{|C_{n}|}\mathcal{MC}_{\mu}\leq\sum_{g\in\mathcal{B}}\#_{g}(n)K_{g}$ (66)
Since the overwriting mismatch cost $\mathcal{MC}(p_{i}n)\leq S(p_{in})$, therefore, the total mismatch cost is upper bounded by,
$\mathcal{MC}(C_{n},p_{in})\leq\sum_{g\in\mathcal{B}}\#_{g}(n)K_{g}+S(p_{in})$ (67)
We can also express the summation term as
$\sum_{g\in\mathcal{B}}\#_{g}(n)K_{g}=|C_{n}|\sum_{g\in\mathcal{B}}\gamma_{g}(n)K_{g}$ (68)
where $\gamma_{g}(n)=\#_{g}(n)/|C_{n}|$. Note that if $\gamma_{g}(n)$ is a constant, i.e., $\gamma_{g}(n)=\gamma_{g}$ for all $n\in\mathbb{N}$ (or for $n\to\infty$), then we get the upper bound similar to Th. 1:
$\mathcal{MC}(C_{n})\leq|C_{n}|K^{\prime}+S(p_{in})$ (69)
where $K^{\prime}=\sum_{g\in\mathcal{B}}\gamma_{g}K_{g}$.
@@ B.3 Circuit as a periodic process
indicating that the counter is precisely in state $\mu$. Note that $p_{X|Z=\mu}(\boldsymbol{x})$ corresponds to $p^{\mu}(\boldsymbol{x})$, consistent with the notation used throughout the paper and as summarized in Table 1.
@@
which coincides with Eq. 23.

### v3

@@ Appendix B Proofs
From Eq. (63), we have:
$\mathcal{MC}(C_{n},p_{\mathrm{in}})=D(p_{\mathrm{in}}||q_{\mathrm{in}})+\sum_{\mu\in V_{\mathrm{nin}}}\left(\mathrm{I}_{\mu+1}(\mathrm{X}_{\mu};\mathrm{X}_{\mathrm{pa}(\mu)})+\Delta_{\mu}\right),$ (64)
@@
$\Delta_{\mu}:=D\left(p_{\mu}p_{\mathrm{pa}(\mu)}||q_{\mu,\mathrm{pa}(\mu)}\right)-D\left(p_{\mathrm{pa}(\mu)}||q_{\mathrm{pa}(\mu)}\right).$ (65)
We assume that the prior distribution $q_{\mu,\mathrm{pa}(\mu)}$ is a product distribution, i.e,
$q_{\mu,\mathrm{pa}(\mu)}(\boldsymbol{x}_{\mu},\boldsymbol{x}_{\mathrm{pa}(\mu)})=q_{\mu}(\boldsymbol{x}_{\mu})q_{\mathrm{pa}(\mu)}(\boldsymbol{x}_{\mathrm{pa}(\mu)}),$ (66)
for every joint state $(\boldsymbol{x}_{\mu},\boldsymbol{x}_{\mathrm{pa}(\mu)})$. Then, $\Delta_{\mu}$ simplifies as follows:
$\Delta_{\mu}$ $=D\left(p_{\mu}p_{\mathrm{pa}(\mu)}||q_{\mu}q_{\mathrm{pa}(\mu)}\right)-D\left(p_{\mathrm{pa}(\mu)}||q_{\mathrm{pa}(\mu)}\right)$ (67)
$=D\left(p_{\mu}||q_{\mu}\right)+D\left(p_{\mathrm{pa}(\mu)}||q_{\mathrm{pa}(\mu)}\right)-D\left(p_{\mathrm{pa}(\mu)}||q_{\mathrm{pa}(\mu)}\right)$ (68)
$=D\left(p_{\mu}||q_{\mu}\right)$ (69)
We also have
$\mathrm{I}_{\mu+1}(\mathrm{X}_{\mu};\mathrm{X}_{\mathrm{pa}(\mu)})=S(\mathrm{X}_{\mu})-S_{\mu+1}(\mathrm{X}_{\mu}\mid\mathrm{X}_{\mathrm{pa}(\mu)}),$ (70)
where $S(\mathrm{X}_{\mu})=-\sum_{\boldsymbol{x}_{\mu}}p_{\mu}(\boldsymbol{x}_{\mu})\log p_{\mu}(\boldsymbol{x}_{\mu})$, and $S_{\mu+1}(\mathrm{X}_{\mu}\mid\mathrm{X}_{\mathrm{pa}(\mu)})$ is the conditional entropy of the state of gate $\mu$ given the states of its parents $\mathrm{pa}(\mu)$ after the gate has been run. For deterministic gates, $S_{\mu+1}(\mathrm{X}_{\mu}\mid\mathrm{X}_{\mathrm{pa}(\mu)})=0$. More generally, since conditional entropy is non-negative, we have the bound
$\mathrm{I}_{\mu+1}(\mathrm{X}_{\mu};\mathrm{X}_{\mathrm{pa}(\mu)})\leq S(\mathrm{X}_{\mu}).$ (71)
Moreover, the KL divergence can be written as
$D\left(p_{\mu}||q_{\mu}\right)=\mathcal{C}~(p_{\mu}||q_{\mu})-S(\mathrm{X}_{\mu}),$ (72)
where $\mathcal{C}~(p_{\mu}||q_{\mu})$ denotes the cross-entropy of $p_{\mu}$ relative to $q_{\mu}$.
Combining Eqs. (71) and (72), we obtain the upper bound
$\sum_{\mu\in V_{\mathrm{nin}}}\left(\mathrm{I}_{\mu+1}(\mathrm{X}_{\mu};\mathrm{X}_{\mathrm{pa}(\mu)})+\Delta_{\mu}\right)\leq\sum_{\mu\in V_{\mathrm{nin}}}\mathcal{C}~(p_{\mu}||q_{\mu}).$ (73)
Furthermore, for any distribution $p_{\mu}$, the cross-entropy satisfies
$\mathcal{C}~(p_{\mu}||q_{\mu})\leq\log\left(\frac{1}{q_{\mu}^{\min}}\right),$ (74)
where
$q_{\mu}^{\min}:=\min_{x}q_{\mu}(x).$ (75)
If all gates share the same prior, i.e., $q_{\mu}=q_{g}$ for all $\mu$ (subscript $g$ implies non-input gate), this yields
$\sum_{\mu\in V_{\mathrm{nin}}}\left(\mathrm{I}_{\mu+1}(\mathrm{X}_{\mu};\mathrm{X}_{\mathrm{pa}(\mu)})+\Delta_{\mu}\right)\leq|C_{n}|\log\left(\frac{1}{q_{g}^{\min}}\right),$ (76)
where $|C_{n}|$ denotes the number of gates in the circuit.
To obtain an explicit upper bound on the KL divergence $D(p_{\mathrm{in}}||q_{\mathrm{in}})$ in terms of the input length $n$, we assume that the prior distribution over the joint state of the input nodes factorizes across input bits:
$q_{\mathrm{in}}(\boldsymbol{x}_{\mathrm{in}})=q_{\mathrm{in}}(x_{\mathrm{in},1},\ldots,x_{\mathrm{in},n})=\prod_{k=1}^{n}q_{\mathrm{in},k}(x_{\mathrm{in},k}).$ (77)
Moreover, we assume that all marginals are identical, i.e., $q_{\mathrm{in},k}=q_{i}$ for all $k$ (subscript $i$ implies input nodes).
Then, using sub-additivity of KL divergence under product priors, we have
$D(p_{\mathrm{in}}||q_{\mathrm{in}})\leq\sum_{k=1}^{n}D(p_{k}||q_{i}),$ (78)
where $p_{k}$ denotes the marginal of $p_{\mathrm{in}}$ on the $k$-th input node. Since each term is bounded by
$D(p_{k}||q_{i})\leq\log\left(\frac{1}{q^{\min}}\right),$ (79)
with $q^{\min}_{i}:=\min_{x}q_{i}(x)$, it follows that
$D(p_{\mathrm{in}}||q_{\mathrm{in}})\leq n\log\left(\frac{1}{q_{i}^{\min}}\right)=nK_{\mathrm{in}},$ (80)
where $K_{\mathrm{in}}:=\log\left(\frac{1}{q_{i}^{\min}}\right)$. This bound is worst-case over all input distributions $p_{\mathrm{in}}$ and depends only on the minimal support of the prior $q$.
Combining Eq. (76) and (80), and substituting them in (64), we obtain the worst-case upper bound over all input distributions $p_{\mathrm{in}}$ on the total mismatch cost of a circuit:
$\mathcal{MC}(C_{n})\leq|C_{n}|K+nK_{\mathrm{in}}.$ (81)
where $K=\log\left(\frac{1}{q_{g}^{\min}}\right)$ and $K_{\mathrm{in}}=\log\left(\frac{1}{q_{\mathrm{in}}^{\min}}\right)$.
@@
To derive the bound (81), we assumed that all the gates in the circuit have the same prior distribution. But this assumption can be dropped by allowing different logic gates to have different priors. Let $\mathcal{B}$ be a basis of logic gates, and let $q_{g}$ denote the prior distribution associated with each gate type $g\in\mathcal{B}$. Let $_{g}(n)$ be the number of gates of type $g$ appearing in a circuit $C_{n}$ of size $n$. (Here, the size of $C_{n}$ counts only non-input logic gates.) Define
$K_{g}:=\log\left(\frac{1}{q_{g}^{\min}}\right),$ (82)
where $q_{g}^{\min}:=\min_{x}q_{g}(x)$.
By Eqs. (71), (72), and (74), each term in the sum on the right-hand side of Eq. (64) is upper bounded. In particular, if gate $\mu$ is of type $g$, then
$\mathrm{I}_{\mu+1}(\mathrm{X}_{\mu};\mathrm{X}_{\mathrm{pa}(\mu)})+\Delta_{\mu}\leq K_{g}.$ (83)
Therefore, if there are $_{g}(n)$ gates of type $g$ in the circuit, we obtain
$\sum_{\mu\in V_{\mathrm{nin}}}\left(\mathrm{I}_{\mu+1}(\mathrm{X}_{\mu};\mathrm{X}_{\mathrm{pa}(\mu)})+\Delta_{\mu}\right)\leq\sum_{g\in\mathcal{B}}\#_{g}(n)K_{g}.$ (84)
The total mismatch cost is bounded by
$\mathcal{MC}(C_{n},p_{\mathrm{\mathrm{in}}})\leq\sum_{g\in\mathcal{B}}\#_{g}(n)K_{g}+nK_{\mathrm{in}}.$ (85)
We can equivalently write the summation term as
$\sum_{g\in\mathcal{B}}\#_{g}(n),K_{g}=|C_{n}|\sum_{g\in\mathcal{B}}\gamma_{g}(n)K_{g},$ (86)
where $\gamma_{g}(n):=\#_{g}(n)/|C_{n}|$ is the fraction of gates of type $g$ in $C_{n}$.
If $\gamma_{g}(n)$ converges to a constant $\gamma_{g}$ as $n$ grows large (or if $\gamma_{g}$ is a constant for all $n$), then we recover an upper bound analogous to Theorem (1):
$\mathcal{MC}(C_{n})\leq|C_{n}|K^{\prime}+nK_{\mathrm{in}},$ (87)
where
$K^{\prime}:=\sum_{g\in\mathcal{B}}\gamma_{g},K_{g}.$ (88)
@@ B.3 Circuit as a periodic process
indicating that the counter is precisely in state $\mu$. Note that $p_{X|Z=\mu}(\boldsymbol{x})$ corresponds to $p^{\mu}(\boldsymbol{x})$, consistent with the notation used throughout the paper and as summarized in Table (1).
@@
which coincides with Eq. (22).
B.4 Further analysis of mutual information terms
Another way to factor out eq (49)
[…]
By applying the chain rule for mutual information we get:
$\Delta I_{\mu}$ $=\Big[I_{\mu}(X_{\mathrm{pa}(\mu)};X_{\tilde{\mu}})+I_{\mu}(X_{\mu};X_{\tilde{\mu}}\mid X_{\mathrm{pa}(\mu)})\Big]$
$\quad-\Big[I_{\mu+1}(X_{\mathrm{pa}(\mu)};X_{\tilde{\mu}})+I_{\mu+1}(X_{\mu};X_{\tilde{\mu}}\mid X_{\mathrm{pa}(\mu)})\Big].$
Since the parents $X_{\mathrm{pa}(\mu)}$ do not change between $\mu$ and $\mu+1$, the first two terms cancel:
$\Delta I_{\mu}\;=\;I_{\mu}(X_{\mu};X_{\tilde{\mu}}\mid X_{\mathrm{pa}(\mu)})-I_{\mu+1}(X_{\mu};X_{\tilde{\mu}}\mid X_{\mathrm{pa}(\mu)}).$ (98)
For a deterministic gate, after the update $X_{\mu+1}=f(X_{\mathrm{pa}(\mu)})$, the gate output is completely determined by its parents, implying $X_{\mu+1}\!\perp\!\!\!\perp X_{\tilde{\mu}}\mid X_{\mathrm{pa}(\mu)}$, and therefore
$I_{\mu+1}(X_{\mu};X_{\tilde{\mu}}\mid X_{\mathrm{pa}(\mu)})=0,\qquad\Rightarrow\qquad\Delta I_{\mu}\;=\;I_{\mu}(X_{\mu};X_{\tilde{\mu}}\mid X_{\mathrm{pa}(\mu)}).$ (99)
Now decompose $X_{\tilde{\mu}}$ into the circuit’s past and future relative to the update order:
$X_{\tilde{\mu}}=(X_{:\mu},\,X_{\mu:}).$
In a feed-forward (topologically ordered) circuit, $X_{:\mu}$ contains no additional information about $X_{\mu}$ once $X_{\mathrm{pa}(\mu)}$ is given: $I_{\mu}(X_{\mu};X_{:\mu}\mid X_{\mathrm{pa}(\mu)})=0.$ Hence,
$\Delta I_{\mu}$ $=I_{\mu}(X_{\mu};X_{:\mu},X_{\mu:}\mid X_{\mathrm{pa}(\mu)})$
$=I_{\mu}(X_{\mu};X_{:\mu}\mid X_{\mathrm{pa}(\mu)})+I_{\mu}(X_{\mu};X_{\mu:}\mid X_{\mathrm{pa}(\mu)},X_{:\mu})$
$=I_{\mu}(X_{\mu};X_{\mu:}\mid X_{\mathrm{pa}(\mu)}).$ (100)
(When appropriate—for deterministic, acyclic circuits with fixed parents— the conditioning can be dropped, giving $I_{\mu}(X_{\mu};X_{\mu:})$)